i have to use substitution to solve this linear system
x + 4y = 17
5x + 2y = -5

Answer by rfer(16322)

x=-4y+17
-------------
5(-4y+17)+2y=-5
-20y+85+2y=-5
-18y=-90
y=5
---------------
5x+2*5=-5
5x=-15
x=-3
----------------
-3+4*5=17
-3+20=17